### v1

# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

Abstract
Traditional monopoly pricing assumes sellers have full information about consumer valuations. We consider monopoly pricing under limited information, where a seller only knows the mean, variance and support of the valuation distribution. The objective is to maximize expected revenue by selecting the optimal fixed price.
We adopt a distributionally robust framework, where the seller considers all valuation distributions that comply with the limited information. We formulate a maximin problem which seeks to maximize expected revenue for the worst-case valuation distribution. The minimization problem that identifies the worst-case valuation distribution is solved using primal-dual methods, and in turn leads to an explicitly solvable maximization problem. This yields a closed-form optimal pricing policy and a new fundamental principle prescribing when to use low and high robust prices.
We show that the optimal policy switches from low to high prices when variance becomes sufficiently large, yielding significant performance gains compared with existing robust prices that generally decay with market uncertainty. This presents guidelines for when the seller should switch from targeting mass markets to niche markets. Similar guidelines are obtained for delay-prone services with rational utility-maximizing customers, underlining the universality and wide applicability of the low-high pricing principle.
1 Introduction
Optimizing revenue in contemporary markets represents a pivotal challenge, extensively explored within theoretical pricing models spanning disciplines such as operations research, computer science, and economics. Prevailing in the theoretical pricing literature is the assumption that sellers possess complete knowledge of how the market responds to price variations. This assumption relies on a full understanding of the demand function, serving as a prior in Bayesian analysis, and significantly influencing the derived optimal price. However, the theoretical advantage of assuming full knowledge is compromised by the scarcity or unreliability of market data. A principle known as Wilson’s doctrine Wilson (1987) calls for less detailed yet parsimonious models, and the development of more resilient pricing theories, capable of accommodating imprecise market details Carroll (2019).
Robust pricing theories seek strategies that perform well across diverse markets, reducing sensitivity to detailed market knowledge. The interplay between a seller’s pricing strategy and market knowledge then becomes crucial. A complete lack of knowledge may result in an overly pessimistic worst-case scenario, rendering a robust price excessively conservative and impractical. A more realistic approach is to assume partial knowledge. This paper addresses this scenario in the context of a monopolist selling a single type of good to a diverse, independent consumer base.
The monopolist aims to determine the revenue-maximizing price by balancing aggressive (low) pricing for a larger market share against higher pricing for a smaller market share. Each consumer’s valuation of the good is an independent draw from a given valuation distribution with the purchase occurring when this valuation exceeds the price. The expected revenue is defined as the product of the price times the probability that the valuation exceeds the price. While the optimal price is easily determined when the valuation distribution is known Riley and Zeckhauser (1983), Myerson (1981), the challenge arises when only partial market knowledge is available.
This study focuses on the case where only the mean, variance, and an upper limit of the valuation distribution are known. Mean-variance information proves suitable for robust decision-making, capturing both the anticipated future scenario (mean) and the risk associated with less likely scenarios (variance). This choice is motivated by the often scarce or non-existent data faced by sellers, where mean, variance, and upper bounds can be based on expert opinions—intuitive statistics that are easier to estimate than the entire valuation distribution; see e.g. Scarf (1958), Azar and Micali (2012).
The partially-informed seller needs to decide on the price in the face of uncertain market demand, much in the same way as an investor who only knows the risk and return of assets. To do so, the seller employs a maximin decision framework. The maximin price, or robust price, is determined by first establishing the worst-case valuation distribution that minimizes expected revenue for all prices and then maximizing expected revenue by setting the optimal price. This involves solving two optimization problems—a semi-infinite linear program and finding the global optimum of a bimodal function that depends delicately on the price and information parameters. The comprehensive nature of this approach reveals the intricate interplay between partial knowledge and robust pricing, surpassing earlier studies Azar and Micali (2012), Kos and Messner (2015), Suzdaltsev (2018), Chen et al. (2022a, 2023b) and highlighting a novel robust pricing principle emerging only when knowledge includes both the upper bound and variance.
Before we explain this general principle, we consider classical monopoly pricing with a fully informed monopolist who knows that valuations follow some given cumulative distribution function $F$ with probability density function $f$. For any given price $p$, the expected revenue is $p(1-F(p))$ and the monopolist thus solves the maximization problem
$\sup_{p}p\left(1-F(p)\right).$ (1)
First-order conditions then prescribe that the optimal price $p^{*}$ solves
$p^{*}=\frac{1-F(p^{*})}{f(p^{*})}.$ (2)
Figure LABEL:fig:beta shows this optimal price for valuations that follow a Beta distribution on the unit interval. Observe that the optimal price is $\mu$ for zero variance, a feature that holds irrespective of the specific valuation distribution. Another feature that can be observed which provided a strong motivation for our work is the following: Viewed as function of $\sigma$, the optimal price $p^{*}$ first decreases to values below the mean and then increases to values above the mean and eventually to the maximal valuation.
The “U-shaped” nature (or quasi-convexity) of revenue as function of demand dispersion – as displayed in the right panel of Figure LABEL:fig:beta – is a known phenomenon in monopolistic markets Johnson and Myatt (2006), Sun (2012). Rigorous underpinning is provided in Johnson and Myatt (2006) by an explanatory analysis that shows the U-shape occurs naturally by imposing only mild conditions on the valuation distribution. Consequently, when there is sufficiently low dispersion, indicative of relatively uniform consumer valuations, sellers are inclined to cater to “mass markets” through competitive pricing strategies. Conversely, as dispersion increases, the average consumer’s willingness to pay diminishes. The heightened heterogeneity among consumers prompts firms to adopt a strategy of restricting sales to a more specialized “niche market” characterized by higher prices. We add a theory aligned with the above intuition, showing that for a rich class of potential consumer valuations, a robust seller indeed switches from low pricing with large market share to high pricing and small market share in response to increased consumer valuation dispersion.
To understand the main pricing principle in this paper, now assume that the seller only knows the mean, standard deviation and support, and chooses to fit to these summary statistics a Beta distribution. Since the monopolist now only pretends to know the valuation distribution, the optimal price found from assuming a Beta distribution is only an approximation, as the ground truth valuation is likely not to be Beta distributed. In fact, the Beta distribution is only one of infinitely many distributions with the same mean, standard deviation and support, and only serves to calculate the expected revenue. This paper presents an alternative pricing strategy, where the monopolist does not commit to a specific valuation distribution, and instead first determines the worst-case valuation distribution that minimizes demand, and only then solves for the revenue-maximizing price. We prove that this robust maximin price corresponds to the black solid line in Figure LABEL:fig:beta. As this paper’s title says, the monopolist switches from low to high prices when the variance exceeds a certain threshold. Before that threshold, the maximin price decreases with variance, at the threshold the price jumps to a higher value, and continues to increase after the threshold.
Contributions. We reveal a new phenomenon of abruptly switching from low to high pricing when the market volatility becomes sufficiently large. In this way, we contribute to advancing our understanding of posted pricing in the presence of limited market information. Azar and Micali (2012) pioneered this direction for knowledge of the mean and standard deviation, demonstrating that the optimal posted price maximizing expected revenue in the worst case can be expressed as an explicit function that decays with the standard deviation. Our more general results align with theirs, and their pricing function appears prominently in several of our theorems for low-risk markets. Another relevant study by Kos and Messner (2015) considers information about the mean and bound on support, so the seller lacks knowledge about the variance. Kos and Messner (2015) also derive an explicit pricing function which appears in our study when we consider high-risk markets. We summarize the existing works and our new results in Table 1.
Information Pioneering works Pricing insights
Mean Variance Azar and Micali (2012) Carrasco et al. (2018) Azar and Micali’s pioneering work revealed a robust price that decreases with variance. Carrasco et al. considered an upper bound on variance, yielding the same price.
Mean Support Kos and Messner (2015) Kos en Messner found a robust price that increases with the valuation support and is (by definition) insensitive to variance.
Mean Variance Support Suzdaltsev (2018) This paper: -Proposition 4.1 -Proposition 4.2 -Theorem 5.2 This paper and unpublished work of Suzdaltsev present robust prices that first decrease and then increase with variance, causing the switch from low to high pricing strategies in response to increased market volatility. The high pricing strategy targets niche markets.
Table 1: Pioneering papers and contributions in this paper.
From a more theoretical perspective, we rigorously establish several propositions and theorems (see Table 1) for both the classical pricing problem and pricing in an extended setting with delay-prone services. These theoretical results together provide a firm underpinning of this principle of low and high robust pricing. In all cases, we derive the optimal posted prices for when the seller knows the mean, bound on support, and standard deviation range of the valuation distributions, generalizing all earlier works including Azar and Micali (2012) and Kos and Messner (2015). We obtain explicit expressions for the optimal prices by solving a maximin problem with distributionally robust optimization methods. We solve the minimization problem with a classical primal-dual method, see e.g. Rogosinski (1958), Shapiro (2001), and show that the solution renders a bimodal revenue function with up to three candidate prices that achieve local maxima. Solving the maximization problem then requires showing that the global optimum is always one of these three local optima, ruling out boundary points; the robust seller then always chooses one of these three candidate prices, a crucial insight that generalizes all existing contributions listed in Table 1.
Outline. The paper is structured as follows. Section 2 presents an overview of related literature on monopoly pricing and distributionally robust optimization. Section 3 then presents tight tail bounds for the valuation distribution, and Section 4 the first results on robust pricing and the low-high pricing phenomenon. Section 5 contains the most general result for low-high pricing by showing that the partially informed seller always selects one of three closed-form pricing functions. Section 6 treats pricing in delay-prone services and shows that similar robust pricing strategies are optimal. Section 7 concludes and reflects on future work. Appendix A presents the primal-dual proof of the tight-tail bounds and Appendix B contains the proofs of the pricing theorems.
2 Related literature
Our study adds to the stream of research on robust pricing decisions under limited information, where sellers must commit to fixed prices without possessing complete knowledge of market valuation and the corresponding demand function. Our approach focuses on mean-variance (or standard deviation) information, a widely studied information set in applications such as investment theory Natarajan and Teo (2017), option pricing De Schepper and Heijnen (2007), insurances Roos et al. (2022), service operation management van Eekelen et al. (2022), Taylor (2018) and inventory management Scarf (1958), Gallego (1992). In the management science and operations literature, this information setting is commonly referred to as ‘distribution-free’, highlighting that the decision-making under uncertainty does not require detailed or full distributional knowledge.
Distributionally robust optimization. Such distribution-free analysis combined with optimization falls in the more general domain of distributionally robust optimization (DRO), a mathematical framework that addresses uncertainty in optimization problems by considering an ambiguity set of possible probability distributions for the involved random variables, rather than relying on a single assumed distribution. The ambiguity sets considered in this paper are based on moment information, as in the pioneering work of Scarf (1958) on robust inventory management, and later generalized to a powerful method with many applications Delage and Ye (2010), Wiesemann et al. (2014). Applied to monopoly pricing, this moment-based DRO approach helps create more robust and conservative pricing strategies that perform well under a range of scenarios. Instead of making strong assumptions about the specific valuation distribution, DRO considers a set of possible distributions, forming a certain ambiguity set, and resulting in pricing strategies that are robust across different scenarios, without relying on a single, possibly incorrect, assumed distribution. This aligns well with Occam’s razor Chen et al. (2023a) and Wilson’s doctrine Carroll (2019), two principles of parsimony that aim for explanatory models able to capture the essence without making unnecessary assumptions or requiring overly strong parametric assumptions, such as particular features of probability distributions.
Robust monopoly pricing. Azar and Micali’s foundational result for the information set with mean and variance (and no bound on the support) has been rediscovered and rederived in various papers Carrasco et al. (2018), Chen et al. (2022a), Suzdaltsev (2018). The information set consisting of the triplet mean, variance, and bound on support has remained largely unexplored. A notable exception is Suzdaltsev (2018), where this setting was studied as a special single-bidder case of an auction. In fact, one of the results in this paper (Proposition 4.2) was mentioned without proof in an early draft of Suzdaltsev (2018) and was removed from the final published version. We thus focus on maximizing absolute revenue with information about the first two moments, just as in Azar and Micali (2012), Chen et al. (2022a), Suzdaltsev (2018). Alternative information sets that were studied in the context of robust pricing include knowledge of higher moments Carrasco et al. (2018), percentiles Eren and Maglaras (2010), mean absolute deviation Roos and den Hertog (2021), Elmachtoub et al. (2021), mean preserving contraction and spread Condorelli and Szentes (2020), Roesler and Szentes (2017), or knowing that the valuation distribution is within the proximity of a given reference distribution Bergemann and Schlag (2011), Chen et al. (2023b). Instead of maximin expected revenue, alternative objectives studied in the literature include minimax regret Bergemann and Schlag (2008, 2011), Elmachtoub et al. (2021), Chen et al. (2023b) and the competitive or approximation ratio Eren and Maglaras (2010), Giannakopoulos et al. (2023), Elmachtoub et al. (2021), Wang et al. (2024). Let us also mention that our work connects to Bayesian persuasion Kamenica and Gentzkow (2011), Kamenica (2019), a research line that explores the optimal type or level of information for optimizing individual utility or social welfare.
Regularity assumptions. The lion’s share of the pricing literature assumes strong regularity conditions on the valuation distribution that guarantee existence and uniqueness of an optimal price. In particular, assuming that the valuation has a monotone hazard rate, or some closely related condition, can guarantee that the revenue function is concave and has a unique optimizer Ziya et al. (2004), Schweizer and Szech (2019). Such regularity assumptions greatly simplify the mathematical framework, but true valuation distributions are more often irregular than not Morganti (2022), Roughgarden and Schrijvers (2016). Loertscher and Muir (2022) write: “The standard assumption, which is almost universally maintained in economics, is that expected revenue is concave. The typical justification for this assumption, other than it being standard, is that it is deemed an analytic simplification that permits one to focus on the key economic insights without cluttering the analysis with case distinctions and multiplicity of local maxima. We have never seen it justified on the basis of empirical evidence, and we will not impose it.” See Celis et al. (2014) for empirical evidence that the value distribution need not be regular. See also the literature overview in Loertscher and Muir (2022) for more references on non-concave revenue and a proof that whenever two or more separate markets merge, the revenue is no longer concave. Even in robust settings, regularity conditions are often maintained for mathematical tractability, see e.g. Maskin and Riley (1984), Eren and Maglaras (2010), Allouah et al. (2022). For robust pricing as in this paper, one should let go of regularity completely, as the worst-case distribution that plays a pivotal role in the maximin analysis typically fails this assumption. In fact, if you would add to the minimization problem the condition that the worst-case valuation needs to be regular, we would lose the key take-aways and novel economic insights of this paper.
Pricing in service operations. Monopoly pricing in the context of delay-sensitive service systems has been extensively examined. A well-established body of literature, as detailed by Hassin and Haviv (2003), delves into mathematical models that capture customer decision-making regarding joining delay-sensitive services based on individual utility considerations. The pioneering work by Naor (1969) laid the groundwork for this branch of operations management. Naor’s seminal model views the system as an M/M/1 queue and models customer utility as linearly dependent on the price, value of service, and the anticipated waiting time. Naor revealed that rational individuals have a tendency to overcrowd the system and suggested the introduction of a fixed fee for service that optimizes social welfare. The determination of this socially optimal fee involves solving first-order conditions, akin to classical monopoly pricing. Expanding upon the foundational work by Naor, numerous studies investigated pricing and control considerations for a variety of extended queueing scenarios, with often the focus on determining the price that maximizes revenue or welfare; see e.g. Dewan and Mendelson (1990), Chen and Frank (2004), Ha (1998), Taylor (2018), Baron et al. (2023), Chen et al. (2022b). Stidham (1992) made a significant contribution by introducing the concept of random valuation for individual customers and assuming that the arrival rate of customers with a valuation surpassing a specified threshold. In this framework, valuation is often referred to as willingness-to-pay, and as in classical monopoly pricing, is a random variable with some given distribution. Cachon and Feldman (2011), Haviv and Randhawa (2014) build on Stidham’s framework and assume that a rational customer joins when its random valuation exceeds the sum of the price and the expected delay costs, and determined optimal prices when the valuation distribution is assumed to be known. We will perform a distribution-free analysis of this rational queueing model and show that again the optimal pricing strategy switches from low to high pricing when the standard deviation of the valuation crosses a certain threshold.
3 Robust pricing framework
The optimal Bayesian price $p^{*}$ in (2) strongly depends on the (properties of the) hazard rate $f(p)/(1-F(p))$ of some specific valuation distribution. The robust pricing developed in this paper, however, only uses summary statistics and renders a revenue-guarantee accomplished across all possible valuation distributions, without referring to any specific distribution. Therefore, the basic problem for robust monopoly pricing considered in this paper is
$\sup_{p}R(p,\mathcal{P}),$ (3)
where
$R(p,\mathcal{P})\mathrel{\mathop{:}}=\inf_{\mathbb{P}\in\mathcal{P}}p\mathbb{P}(X\geq p)$ (4)
with $\mathbb{P}$ some valuation distribution and $\mathcal{P}$ the ambiguity set with all valuation distributions that comply with the partial knowledge that the seller has. Specifically, we consider the ambiguity set (with $\mathcal{B}$ the Borel sigma-algebra over $\mathbb{R_{+}}$)
$\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)=\left\{\mathbb{P}:\mathcal{B}\to[0,1]\mid\mathbb{P}(X\in[0,\beta])=1,\,\mathbb{E}(X)=\mu,\mathbb{E}(X^{2})\in[\underaccent{\bar}{\sigma}^{2}+\mu^{2},\bar{\sigma}^{2}+\mu^{2}]\right\}$ (5)
containing all distributions with mean $\mu$, standard deviation in the interval $[\underaccent{\bar}{\sigma},\bar{\sigma}]$ and support upper bounded by $\beta$. For this ambiguity set, we determine tight bounds for the tail probabilities, solving the minimization part of the central maximin problem (3):
Lemma 3.1 (Tight tail bounds).
Consider ambiguity set $\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ and denote $\bar{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\bar{\sigma}^{2}}{\beta-\mu}$, $\underaccent{\bar}{\upsilon}_{1}\mathrel{\mathop{:}}=\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{\beta-\mu}$ and $\underaccent{\bar}{\upsilon}_{2}\mathrel{\mathop{:}}=\mu+\frac{\underaccent{\bar}{\sigma}^{2}}{\mu}$. Then,
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p)=\begin{cases}\frac{(\mu-p)^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}},\quad&p\in(0,\bar{\upsilon}_{1}],\\ \frac{\mu-p}{\beta-p},\quad&p\in[\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1}],\\ \frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)},\quad&p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}],\\ 0,\quad&p\in[\underaccent{\bar}{\upsilon}_{2},\beta].\end{cases}$ (6)
Proof (sketch). Lemma 3.1 generalizes the one-sided Chebyshev bound, also known as Cantelli’s inequality, for mean-variance information Feller (1971), and also extends results in De Schepper and Heijnen (1995) for precise knowledge of the variance. Let us sketch the proof of Lemma 3.1. The key step is to formulate the tight bound as the solution to a semi-infinite linear program, see e.g. Shapiro (2001). Writing the tail probability $\mathbb{P}(X>p)$ as the expectation of an indicator function, $\int_{x}\mathds{1}{\{x>p\}}{\rm d}\mathbb{P}(x)$, we see that the tight bound follows from solving the primal problem (with $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$)
$\!\inf_{\mathbb{P}\in\mathcal{M}}$ $\int_{x}\mathds{1}{\{x>p\}}{\rm d}\mathbb{P}(x),$ (7)
s.t. $\int_{x}{\rm d}\mathbb{P}(x)=1,\ \int_{x}x{\rm d}\mathbb{P}(x)=\mu,\ \int_{x}x^{2}{\rm d}\mathbb{P}(x)\geq\underaccent{\bar}{\sigma}^{2}+\mu^{2},\ \int_{x}x^{2}{\rm d}\mathbb{P}(x)\leq\bar{\sigma}^{2}+\mu^{2}.$
The proof then continues with solving the corresponding dual problem
$\sup_{\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}}$ $\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2}),$ (8)
s.t. $\mathds{1}{\{x>p\}}\geq\lambda_{0}+\lambda_{1}x+(\lambda_{2}+\lambda_{3})x^{2}=:F(x)\ \forall x\in[0,\beta],\lambda_{2}\geq 0,\lambda_{3}\leq 0,$
by considering all possible shapes of the minorant $F(x)$. For each shape we solve the dual and use complementary slackness to obtain a primal solution, showing that the primal and dual solutions agree. The full proof is presented in Section A. ∎
The solution to (4) in Lemma 3.1 will be leveraged for solving the remaining maximization problem (3) to obtain robust pricing strategies for various choices of information sets that reflect the knowledge of the partially informed seller. In Section 4 we consider the cases where the seller knows either the precise value of the variance or the upper bound on the support (but not both). Building on the results in section 4, we present optimal robust prices in Section 5 for the most general setting considered in this paper, where the seller has the knowledge captured in the ambiguity set (5). In Section 6 we consider a more advanced pricing model for delay-prone services for which we will again employ Lemma 3.1 to perform the maximin analysis.
4 Robust pricing with two prices
This section develops some first robust pricing insights, focusing on two scenarios: when the seller is aware of either the exact value or the maximum value of the standard deviation. We demonstrate that the optimal robust pricing strategy involves choosing one of two prices, as outlined in Propositions 4.1 and 4.2. These propositions play a pivotal role in Section 5, where we extend our analysis to the broader context of knowing the standard deviation falls within a specified interval. In this case, three candidate prices emerge for the optimal pricing strategy.
4.1 Upper bound on standard deviation
First assume that the seller knows the mean valuation $\mu$, the maximal valuation $\beta$, and that the standard deviation of the valuation distribution is at most $\bar{\sigma}$. The robust price should then solve the maximin problem $\sup_{p}\bar{R}(p)$, where the function $\bar{R}(p)$ follows from Lemma 3.1 for $\underaccent{\bar}{\sigma}=0$:
$\bar{R}(p)\mathrel{\mathop{:}}=R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))=\begin{cases}R_{1}(p)\mathrel{\mathop{:}}=\frac{p(\mu-p)^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}},\quad&p\in(0,\bar{\upsilon}_{1}],\\ R_{2}(p)\mathrel{\mathop{:}}=\frac{p(\mu-p)}{\beta-p},\quad&p\in[\bar{\upsilon}_{1},\mu],\\ 0,\quad&p\in[\mu,\beta],\end{cases}$ (9)
with $\mathcal{P}(\mu,\bar{\sigma},\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and upper bound on standard deviation $\bar{\sigma}$. In this case, the worst-case revenue $\bar{R}(p)$ can thus be viewed as the maximum of $R_{1}(p)$, $R_{2}(p)$ and 00, so the optimal robust price will follow from either $R_{1}(p)$ or $R_{2}(p)$. This is illustrated in Figure LABEL:figmain1. Observe that $\bar{R}(p)$ indeed consists of the two concatenated functions $R_{1}(p)$, $R_{2}(p)$, and that both functions might create a local maximum. For $\bar{\sigma}=0.36$ there are two local maxima, corresponding to two prices, and the global maximum follows from the largest price. As it turns out, the local maximum delivered by $R_{1}(p)$, associated with a relatively low price, is globally optimal when the risk level $\bar{\sigma}$ is sufficiently small. And vice versa, when $\bar{\sigma}$ is sufficiently large, the optimal price is relatively high and corresponds to the argmax of $R_{2}(p)$.
We will now formalize these observations in the first main pricing theorem. To do so, we define the two price functions
$p^{*}_{m}(\beta)$ $=\beta-\sqrt{\beta(\beta-\mu)},$ (10)
and
$p^{*}_{l}(\bar{\sigma})$ $=\mu-\bar{\sigma}\kappa(\mu,\bar{\sigma}),$ (11)
with
$\kappa(\mu,\bar{\sigma})$ $\mathrel{\mathop{:}}=\left(\tfrac{\mu}{\bar{\sigma}}+\sqrt{1+(\tfrac{\mu}{\bar{\sigma}})^{2}}\right)^{\frac{1}{3}}+\left(\tfrac{\mu}{\bar{\sigma}}-\sqrt{1+(\tfrac{\mu}{\bar{\sigma}})^{2}}\right)^{\frac{1}{3}}.$ (12)
The arguments of these price functions are chosen deliberately to express the crucial fact that the price $p^{*}_{l}(\bar{\sigma})$ responds to the risk level $\bar{\sigma}$, while the price $p^{*}_{m}(\beta)$ does not. Here is then our first pricing result, which shows that for a large enough maximal risk-level $\bar{\sigma}$, the optimal price becomes risk-insensitive:
Proposition 4.1 (Risk-sensitive and risk-insensitive pricing).
The optimal robust price $\bar{p}^{*}=\arg\sup_{p}\bar{R}(p)$ with $\bar{R}(p)$ in (9) is
$\bar{p}^{*}=\begin{cases}p^{*}_{l}(\bar{\sigma})\quad&{\rm if}\ \bar{\sigma}\leq\bar{\sigma}^{*},\\ p^{*}_{m}(\beta)\quad&{\rm if}\ \bar{\sigma}\geq\bar{\sigma}^{*},\end{cases}$ (13)
with $\bar{\sigma}^{*}=\left(\tfrac{32}{27}(\beta-\mu)\left(\sqrt{\beta(\beta-\mu)}-(\beta-\mu)\right)\right)^{\frac{1}{2}}$.
Proof. The derivative of $R_{1}(p)$ equals zero at $p^{*}_{l}(\bar{\sigma})$ and $\mu$, but only when $p^{*}_{l}(\bar{\sigma})$ lies in the interval $(0,\bar{\upsilon}_{1}]$ is this the maximum ($\mu$ is never the maximum). Since $R_{1}(p)$ increases in $[0,p^{*}_{l}(\bar{\sigma})]$, the maximum is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\min\{p^{*}_{l}(\bar{\sigma}),\bar{\upsilon}_{1}\}$. Similarly, $R_{2}(p)$ is concave and attains a unique maximum at $p^{*}_{m}(\beta)\in[0,\mu]$. Since $R_{2}(p)$ increases in $[0,p^{*}_{m}(\beta)]$, the maximum of $R_{2}(p)$ in $[\bar{\upsilon}_{1},\mu]$ is attained in $\bar{\upsilon}_{1}$ when $\bar{\upsilon}_{1}=\max\{p^{*}_{m}(\beta),\bar{\upsilon}_{1}\}$. We next show that, despite $\bar{\upsilon}_{1}$ being a candidate local maximizer, it can never be the global maximizer for the function $\bar{R}(p)$. Observe that
$R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))$ $=R(p,\mathcal{P}(\mu,\bar{\sigma})\cap\mathcal{P}(\mu,\beta))\geq\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\},$ (14)
and
$R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))$ $=R(p,\mathcal{P}(\mu,\bar{\sigma}))\mathds{1}\{p\in(0,\bar{\upsilon}_{1}]\}+R(p,\mathcal{P}(\mu,\beta))\mathds{1}\{p\in(\bar{\upsilon}_{1},\beta]\}$
$\leq\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\}.$ (15)
Hence, $R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))=\max\{R(p,\mathcal{P}(\mu,\sigma)),R(p,\mathcal{P}(\mu,\beta))\}$. Consequently, since the maximization problem is equivalent to the maximization problem without either $\bar{\sigma}$ or $\beta$, we have $\bar{p}^{*}\in\{p^{*}_{l}(\bar{\sigma}),p^{*}_{m}(\beta)\}$. Since $R_{1}$ is continuously differentiable and attains its maximum in $p^{*}_{l}(\bar{\sigma})$, we can employ the envelope theorem to show that
$\frac{\partial R_{1}(p^{*}_{l}(\bar{\sigma}))}{\partial\bar{\sigma}^{2}}=\frac{\partial R_{1}(p)}{\partial\bar{\sigma}^{2}}\biggr{\rvert}_{p=p^{*}_{l}(\bar{\sigma})}=-\frac{p^{*}_{l}(\bar{\sigma})(\mu-p^{*}_{l}(\bar{\sigma}))^{2}}{((\mu-p^{*}_{l}(\bar{\sigma}))^{2}+\bar{\sigma}^{2})^{2}}\leq 0,$ (16)
since $p^{*}_{l}(\bar{\sigma})\geq 0$. Hence, $R_{1}(p^{*}_{l}(\bar{\sigma}))$ is non-increasing in $\bar{\sigma}$, so that $\bar{\sigma}^{*}$ is the unique value of $\bar{\sigma}$ that satisfies $R_{1}(p^{*}_{l}(\bar{\sigma}))=R_{2}(p^{*}_{m}(\beta))$, which proves the claim. ∎
One important conclusion from Proposition 4.1 is that pricing becomes risk-insensitive when the upper bound on variance is sufficiently large; see Figure LABEL:figmain1. However, this will change drastically when the seller has better or more precise variance information.
4.2 Precise knowledge of standard deviation
We will thus consider the setting where the seller knows the precise standard deviation which means $\underaccent{\bar}{\sigma}=\bar{\sigma}$. For notational convenience, we will drop the bar notation and write $\sigma$ instead. We solve
$\sup_{p}\hat{R}(p),$ (17)
where it follows from Lemma 3.1 for $\underaccent{\bar}{\sigma}=\bar{\sigma}$ that
$\hat{R}(p)\mathrel{\mathop{:}}=R(p,\mathcal{P}(\mu,\sigma,\beta))=\begin{cases}R_{1}(p)=\frac{p(\mu-p)^{2}}{(\mu-p)^{2}+\sigma^{2}},\quad&p\in(0,\underaccent{\bar}{\upsilon}_{1}],\\ R_{3}(p)\mathrel{\mathop{:}}=\frac{p(\mu(\mu-p)+\sigma^{2})}{\beta(\beta-p)},\quad&p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}],\\ 0,\quad&p\in[\underaccent{\bar}{\upsilon}_{2},\beta],\end{cases}$ (18)
with $\mathcal{P}(\mu,\sigma,\beta)$ the set of all valuation distribution with mean $\mu$, upper bound on valuation $\beta$, and standard deviation $\sigma$. Figure LABEL:figmain1h displays the worst-case revenue for several risk levels. Observe that $\hat{R}(p)$ still consists of the two quadratic functions, pasted together, and that both functions present a candidate for the global maximum. However, compared with Figure LABEL:figmain1, there are some notable differences. First, the price range with non-degenerate prices increases with $\sigma$, now allowing robust prices that exceed the mean $\mu$. Second, the risk level from which the high price starts dominating the low price seems to be lower than in Figure LABEL:figmain1. And third, the plots in Figure LABEL:figmain1h show that for high prices, the revenue increases significantly with $\sigma$, which leads to premium pricing targeted at users with the highest valuations. We will substantiate these observations in a second pricing theorem, for which we introduce a third price function
$p^{*}_{h}(\sigma,\beta)$ $=\beta-\sqrt{\beta\left(\beta-\mu-\frac{\sigma^{2}}{\mu}\right)},$ (19)
which unlike $p^{*}_{l}(\sigma)$ and $p^{*}_{m}(\beta)$ depends on both $\sigma$ and $\beta$. This price function turns out pivotal in high-risk settings:
Proposition 4.2 (Low and high pricing).
The optimal robust price $\hat{p}^{*}=\arg\sup_{p}{\hat{R}}(p)$ with ${\hat{R}}(p)$ in (18) is
$\hat{p}^{*}=\begin{cases}p^{*}_{l}(\sigma)\quad&{\rm if}\ \sigma\leq\sigma^{*},\\ p^{*}_{h}(\sigma,\beta)\quad&{\rm if}\ \sigma\geq\sigma^{*},\end{cases}$ (20)
with $\sigma^{*}$ defined implicitly as the solution to $R_{1}(p^{*}_{l}(\sigma))=R_{3}(p^{*}_{h}(\sigma,\beta))$.
Figure LABEL:figmain1hopt illustrates Proposition 4.2. In response to increased risk, the robust seller initially lowers the price, aligned with classical pricing theory. When the risk surpasses some threshold, however, the seller changes its strategy drastically and switches to a high price that increases as a function of the risk. This high price function typically exceeds the mean valuation and converges to the maximal valuation when the risk approaches its maximal value.
Proving Proposition 4.2 is considerably more challenging than proving Proposition 4.1. Our proof starts with the following result:
Lemma 4.1.
Consider $\hat{p}^{*}=\arg\sup_{p}{\hat{R}}(p)$ with ${\hat{R}}(p)$ in (18). The optimal price $\hat{p}^{*}$ is $p^{*}_{l}(\sigma)$, $p^{*}_{h}(\sigma,\beta)$ or $\underaccent{\bar}{\upsilon}_{1}$.
Proof. From Proposition 4.1 we know that the maximum of $R_{1}(p)$ is either attained in $p^{*}_{l}(\sigma)$ or in $\underaccent{\bar}{\upsilon}_{1}$. Also, as $\sigma^{2}$ is bounded from above by $\mu(\beta-\mu)$,
$R_{3}^{\prime\prime}(p)=-\frac{2(\mu(\beta-\mu)-\sigma^{2})}{(\beta-p)^{3}}\leq 0.$ (21)
Hence, $R_{3}(p)$ is concave and attains a unique maximum at $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Since $R_{3}(p)$ decreases in $[p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{2}]$, the maximum of $R_{3}(p)$ in $[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$ is attained in $\underaccent{\bar}{\upsilon}_{1}$ when $\underaccent{\bar}{\upsilon}_{1}=\max\{p^{*}_{h}(\sigma,\beta),\underaccent{\bar}{\upsilon}_{1}\}$. ∎
As in the proof Proposition 4.1, we next want to rule out $\underaccent{\bar}{\upsilon}_{1}$ as candidate optimal price, but this now turns out more challenging. The next lemma states that indeed $\underaccent{\bar}{\upsilon}_{1}$ can never be the optimal price, and the proof using some advanced analysis is deferred to the appendix.
Lemma 4.2.
Consider $\hat{p}^{*}=\arg\sup_{p}{\hat{R}}(p)$ with ${\hat{R}}(p)$ in (18). The optimal price $\hat{p}^{*}$ can never be $\underaccent{\bar}{\upsilon}_{1}$.
Lemmas 4.1 and 4.2 together show that the optimal price is either $p^{*}_{l}(\sigma)$ or $p^{*}_{h}(\sigma,\beta)$. The remaining step is to show that the low (high) price is optimal for low (high) variance.
Lemma 4.3.
Consider $\sigma^{*}$ defined in Proposition 4.2. Then, $\hat{R}(p^{*}_{l}(\sigma))>\hat{R}(p^{*}_{h}(\sigma,\beta))$ for $\sigma<\sigma^{*}$, $\hat{R}(p^{*}_{l}(\sigma))<\hat{R}(p^{*}_{h}(\sigma,\beta))$ for $\sigma>\sigma^{*}$ and $\hat{R}(p^{*}_{l}(\sigma))=\hat{R}(p^{*}_{h}(\sigma,\beta))$ for $\sigma=\sigma^{*}$.
Proof. Since $R_{1}(p)$ is maximal and continuously differentiable in $p^{*}_{l}(\sigma)$, we can employ the envelope theorem to show that
$\frac{\partial R_{1}(p^{*}_{l}(\sigma))}{\partial\sigma^{2}}=\frac{\partial R_{1}(p)}{\partial\sigma^{2}}\biggr{\rvert}_{p=p^{*}_{l}(\sigma)}=-\frac{p^{*}_{l}(\sigma)(\mu-p^{*}_{l}(\sigma))^{2}}{((\mu-p^{*}_{l}(\sigma))^{2}+\sigma^{2})^{2}}\leq 0,$ (22)
since $p^{*}_{l}(\sigma)\geq 0$. Similarly, since $R_{3}(p)$ is maximal and continuously differentiable in $p^{*}_{h}(\sigma,\beta)$, the envelope theorem yields
$\frac{\partial R_{3}(p^{*}_{h}(\sigma,\beta))}{\partial\sigma^{2}}=\frac{\partial R_{3}(p)}{\partial\sigma^{2}}\biggr{\rvert}_{p=p^{*}_{h}(\sigma,\beta)}=\frac{p^{*}_{h}(\sigma,\beta)}{\beta(\beta-p^{*}_{h}(\sigma,\beta))}\geq 0,$ (23)
since $p^{*}_{h}(\sigma,\beta)\geq 0$. Together these two statements prove that the threshold value $\sigma^{*}$ exists and is unique. Next, consider $\sigma\xrightarrow{}0$ and assume $p^{*}_{h}(\sigma,\beta)\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$. Then $p^{*}_{h}(\sigma,\beta)=\mu$, which is a contradiction, and $\hat{p}^{*}=p^{*}_{l}(\sigma)$. Furthermore, consider $\sigma\xrightarrow{}\sigma_{\textup{max}}=\sqrt{\mu(\beta-\mu)}$ and assume $p^{*}_{l}(\sigma)\in[0,\underaccent{\bar}{\upsilon}_{1}]$. Then $p^{*}_{l}(\sigma)=0$, which cannot be optimal, so $\hat{p}^{*}=p^{*}_{h}(\sigma,\beta)$. ∎
Proposition 4.2 then follows from Lemmas 4.1, 4.2 and 4.3.
5 Robust pricing with three prices
We proceed with the most general setting covered in this study, where instead of knowing an upper bound or exact value, the seller knows that the standard deviation is contained in an interval $[\underaccent{\bar}{\sigma},\bar{\sigma}]$. The width of this interval thus reflects how confident the seller is about the actual value $\sigma$, or equivalently, how accurate the seller can estimate $\sigma$. From Lemma 3.1 we see that the revenue maximization problem becomes $\sup_{p}\tilde{R}(p)$ with
$\tilde{R}(p)\mathrel{\mathop{:}}=R(p,\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta))=\begin{cases}R_{1}(p)=\frac{p(\mu-p)^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}},\quad&p\in(0,\underaccent{\bar}{\upsilon}_{1}],\\ R_{2}(p)=\frac{p(\mu-p)}{\beta-p},\quad&p\in[\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1}],\\ R_{3}(p)=\frac{p(\mu(\mu-p)+\underaccent{\bar}{\sigma}^{2})}{\beta(\beta-p)},\quad&p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}],\\ 0,\quad&p\in[\underaccent{\bar}{\upsilon}_{2},\beta].\end{cases}$ (24)
Here is this paper’s most general and main result for robust monopoly pricing:
Theorem 5.1 (Three prices theorem).
When the seller knows $\mu$, $\beta$, and that $\sigma\in[\underaccent{\bar}{\sigma},\bar{\sigma}]$, the optimal robust price $p^{*}=\arg\sup_{p}\tilde{R}(p)$ is contained in the set $\{p^{*}_{l}(\bar{\sigma}),p^{*}_{m}(\beta),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\}$.
Proof. When $\underaccent{\bar}{\sigma}=\bar{\sigma}$, we know from Proposition 4.2 that $p^{*}\in\{p^{*}_{l}(\bar{\sigma}),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\}$. For the remainder of the proof, assume $\underaccent{\bar}{\sigma}<\bar{\sigma}$. From Proposition 4.1 and 4.2, we know that $p^{*}_{l}(\bar{\sigma}),p^{*}_{m}(\beta),p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ are the optimal prices of the unconstrained functions $R_{1}(p),R_{2}(p),R_{3}(p)$ respectively. What remains is to exclude the potential boundary optimal prices $\bar{\upsilon}_{1}$ and $\underaccent{\bar}{\upsilon}_{1}$.
We will now show that $\bar{\upsilon}_{1}$ cannot be the unique optimal price. Since $\bar{\upsilon}_{1}\in(0,\underaccent{\bar}{\upsilon}_{1})$, we can assume without loss of generality that $p^{*}<\underaccent{\bar}{\upsilon}_{1}$ (since $p^{*}\geq\underaccent{\bar}{\upsilon}_{1}$ implies $p^{*}\neq\bar{\upsilon}_{1}$). The following holds due to Proposition 4.1:
$R(p,\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta))$ $\geq R(p,\mathcal{P}(\mu,\bar{\sigma},\beta))=\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\}.$ (25)
In addition, it holds that
$R(p,\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta))$ $=R(p,\mathcal{P}(\mu,\bar{\sigma}))\mathds{1}\{p\in(0,\bar{\upsilon}_{1}]\}+R(p,\mathcal{P}(\mu,\beta))\mathds{1}\{p\in(\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1})\},$
$\leq\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\}.$ (26)
Hence, $R(p,\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta))=\max\{R(p,\mathcal{P}(\mu,\bar{\sigma})),R(p,\mathcal{P}(\mu,\beta))\}$ and therefore, $p^{*}$ is either $p^{*}_{l}(\bar{\sigma})$ or $p^{*}_{m}(\beta)$ whenever $p^{*}\in[0,\underaccent{\bar}{\upsilon}_{1})$, since the maximization problem is equivalent to the maximization problem without either $\bar{\sigma}$ or $\beta$. This means we can exclude $\bar{\upsilon}_{1}$ as an optimal price candidate.
Next, we will show that $\underaccent{\bar}{\upsilon}_{1}$ cannot be the unique optimal price. Notice that for $\underaccent{\bar}{\upsilon}_{1}$ to be the unique global maximum, the ordering $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)<\underaccent{\bar}{\upsilon}_{1}<p^{*}_{m}(\beta)$ has to hold. However, $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\geq p^{*}_{m}(\beta)$, as both prices exhibit the form of $\beta-\sqrt{\beta(\beta-c)}$ with $c$ being $\mu$ for $p^{*}_{m}(\beta)$ and $\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu}$ for $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$. Hence, we can exclude $\underaccent{\bar}{\upsilon}_{1}$ as an optimal price candidate as well. ∎
Although Theorem 5.1 largely resolves the robust pricing problem, as it reduces the set of infinitely many prices to three prices only, it does not provide insight into how the optimal price responds to changes in market volatility $\sigma$ or market valuation bound $\beta$. We therefore build on Theorem 5.1, and the extreme cases covered in Theorems 4.1 and 4.2, to obtain a more detailed characterization of the optimal pricing strategy.
Theorem 5.1 tells us that the optimal price can be found by three pairwise comparisons, for which we shall introduce three functions. The function $f$ is the value of $\bar{\sigma}$ for which $R_{1}(p^{*}_{l}(\bar{\sigma}))=R_{2}(p^{*}_{m}(\beta))$ and corresponds with the threshold value $\bar{\sigma}$ that appeared in Proposition 4.1:
$f(\mu,\beta)=\left(\tfrac{32}{27}(\beta-\mu)\left(\sqrt{\beta(\beta-\mu)}-(\beta-\mu)\right)\right)^{\frac{1}{2}}.$ (27)
The function $g$ is the value of $\underaccent{\bar}{\sigma}$ for which $R_{2}(p^{*}_{m}(\beta))=R_{3}(p^{*}_{h}(\underaccent{\bar}{\sigma},\beta))$. Since $R_{2}(p^{*}_{m}(\beta))$ does not depend on $\underaccent{\bar}{\sigma}$, we write
$g(\mu,\beta)=\left(2\beta\sqrt{R_{2}(p^{*}_{m}(\beta))\mu}-\beta R_{2}(p^{*}_{m}(\beta))-\mu^{2}\right)^{\frac{1}{2}}.$ (28)
The function $h$ is the value of $\underaccent{\bar}{\sigma}$ for which $R_{1}(p^{*}_{l}(\bar{\sigma}))=R_{3}(p^{*}_{h}(\underaccent{\bar}{\sigma},\beta))$. Since $R_{1}(p^{*}_{l}(\bar{\sigma}))$ does not depend on $\underaccent{\bar}{\sigma}$, we write
$h(\bar{\sigma},\mu,\beta)=\left(2\beta\sqrt{R_{1}(p^{*}_{l}(\bar{\sigma}))\mu}-\beta R_{1}(p^{*}_{l}(\bar{\sigma}))-\mu^{2}\right)^{\frac{1}{2}}.$ (29)
We can now characterize the optimal price in terms of risk.
Theorem 5.2.
The optimal robust price $p^{*}=\arg\sup_{p}\tilde{R}(p)$ is
$p^{*}=\begin{cases}p^{*}_{l}(\bar{\sigma})\quad&{\rm if}\ (\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{l},\\ p^{*}_{m}(\beta)\quad&{\rm if}\ (\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{m},\\ p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)\quad&{\rm if}\ (\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{h},\end{cases}$ (30)
where the risk threshold sets are given by
$\Sigma_{l}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq h(\bar{\sigma},\mu,\beta),\bar{\sigma}\leq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$ (31)
$\Sigma_{m}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\leq g(\mu,\beta),\bar{\sigma}\geq f(\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\},$
$\Sigma_{h}=\{(\underaccent{\bar}{\sigma},\bar{\sigma})\in\mathbf{R}_{+}^{2}|\underaccent{\bar}{\sigma}\geq g(\mu,\beta),\underaccent{\bar}{\sigma}\geq h(\bar{\sigma},\mu,\beta),\underaccent{\bar}{\sigma}\leq\bar{\sigma}\}.$
Figure 4 visualizes Theorem 5.2. When either $\underaccent{\bar}{\sigma}$ or $\bar{\sigma}$ is held constant and does not coincide with the intersection of $\Sigma_{l}$ and $\Sigma_{m}$, varying the other parameter will result in at most two relevant prices, but never three. The high pricing strategy $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ becomes optimal when $\underaccent{\bar}{\sigma}$ and $\bar{\sigma}$ are both sufficiently large.
[…]
Figure 4: Three prices theorem for $\mu=0.5$ and $\beta=1$.
5.1 Comparison with optimal pricing
As is common in the pricing literature, we will now evaluate the performance of the robust pricing strategy with the optimal Bayesian strategy; see e.g. Azar and Micali (2012) and Chen et al. (2022a). The optimal Bayesian strategy is known to be fixed pricing, see Riley and Zeckhauser (1983), and achieves the optimal expected revenue
$\textup{OPT}(\mathbb{P})\mathrel{\mathop{:}}=\sup_{p}p\mathbb{P}(X\geq p).$ (32)
Markov’s inequality yields
$\mathbb{P}(X\geq p)\leq\frac{\mu}{p}\Leftrightarrow p\mathbb{P}(X\geq p)\leq\mu,$ (33)
and hence $\textup{OPT}(\mathbb{P})\leq\mu$. This is in fact a tight bound. To see this, let $p\in[\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu},\mu+{\bar{\sigma}^{2}}/{\mu}]$ and consider the two-point distribution $\mathbb{P}_{0}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ with atoms $\{0,p\}$ and associated probabilities $1-\mu/p$ and ${\mu}/{p}$, respectively, for which $p\mathbb{P}_{0}(X\geq p)=\mu$. Combining this bound with Theorem 5.2 yields, upon some rewriting, the following performance guarantee for robust pricing:
Corollary 5.1.
For any $\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ consider the optimal price $p^{*}$ in (30). Then,
$\frac{p^{*}\mathbb{P}(X\geq p^{*})}{\textup{OPT}(\mathbb{P})}\geq\begin{cases}\rho_{l}(\mu,\bar{\sigma})\mathrel{\mathop{:}}=\frac{1}{2}\frac{\bar{\sigma}}{\mu}\kappa^{3}(\mu,\bar{\sigma})\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{l},\\ \rho_{m}(\mu,\beta)\mathrel{\mathop{:}}=\frac{2}{\mu}(\beta-\sqrt{\beta(\beta-\mu)})-1\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{m},\\ \rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)\mathrel{\mathop{:}}=2(1-\sqrt{1-\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}})-\tfrac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}\quad&\text{\rm if }(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{h}.\end{cases}$ (34)
[…]
Figure 5: Corollary 5.1 with $\mu=0.5$ and $\beta=1$.
Figure 5 illustrates Corollary 5.1. When the seller has precise information about the standard deviation, there occurs a U-shaped performance guarantee, aligned with the U-shaped revenue curves found in monopolistic markets Johnson and Myatt (2006), Sun (2012) discussed in the introduction. In our case, this U-shape occurs as the seller switches from low to high pricing, changing the curve from being decreasing to increasing. More generally, we have found that the U-shape occurs when the seller has precise enough (not perfect per se) information on the standard deviation. When the seller only knows an upper bound, the U-shape is replaced with a function that first decreases and then remains constant, see Figure 5. We substantiate these numerical insights with the following result:
Corollary 5.2 (Properties of price and revenue).
(i) $p^{*}_{l}(\bar{\sigma})$ and $\rho_{l}(\mu,\bar{\sigma})$ increase in $\mu$ and decrease in $\bar{\sigma}$; (ii) $p^{*}_{m}(\beta)$ and $\rho_{m}(\mu,\beta)$ increase in $\mu$ and decrease in $\beta$; (iii) $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ and $\rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)$ increase in $\underaccent{\bar}{\sigma}$ and decrease in $\beta$.
The optimal price and performance guarantee thus both move in the same direction in response to changing mean, support and variance. In volatile markets, the essence of successful pricing lies in its ability to adjust prices in accordance with uncertainty levels, ensuring higher prices with increasing volatility. Corollary 5.2 shows that out of the three potential optimal price functions, only the high price is positively correlated with variance, resulting in a performance guarantee that also increases with variance.
5.2 Effect of bundling
The existing literature considering mean-variance information described a pricing policy with only one robust price Azar and Micali (2012), Chen et al. (2022a), the low price in our pricing policy. We also include support information, and find two and at most three price functions. Our new pricing policy that adjusts the price to market volatility is transparent and simple with as main strength a high price responding to high uncertainty, leading to high revenue. If the goal is to use only one, instead of two or three prices, the seller could opt for bundling. The seller then sets a price for a bundle of goods, so that customers buy either the bundle or nothing. A big theme in the pricing literature centers around the question whether bundling is better than separate sales, and in general there is no clear or straightforward answer to that question Bakos and Brynjolfsson (1999), Hart and Nisan (2017). We will use the result for the single product to assess this bundle product pricing. Assuming valuations are i.i.d., the revenue of selling a bundle of $m$ goods with valuations $X_{1},\ldots,X_{m}$ for a bundle price $p_{(m)}$ has expected revenue
$p_{(m)}\cdot\mathbb{P}\Big{(}\sum_{i=1}^{m}X_{i}\geq p_{(m)}\Big{)}.$ (35)
Turning to the limited information setting, the single-product information about the mean, variance and maximal valuation translates into information for the bundle by multiplying all three characteristics with the number of goods $m$. Notice that a bundle of $m$ i.i.d. goods, each with mean $\mu$, standard deviation $\sigma\in[\underaccent{\bar}{\sigma},\bar{\sigma}]$ and maximal valuation $\beta$ can be viewed as selling a single good with as valuation the average of all $m$ valuations, so with mean $\mu$, maximal value $\beta$ and $\sigma\in[\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\frac{\bar{\sigma}}{\sqrt{m}}]$. As a result, for any $p<\mu$ and large enough $m$, the tail bound (6) becomes
$\inf_{\mathbb{P}\in\mathcal{P}\left(\mu,\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\frac{\bar{\sigma}}{\sqrt{m}},\beta\right)}\mathbb{P}(X\geq p)=\frac{(\mu-p)^{2}}{(\mu-p)^{2}+\tfrac{\bar{\sigma}^{2}}{m}}$ (36)
and the expected worst-case revenue is maximized by the low price $p^{*}_{l}(\mu,\frac{\sigma}{\sqrt{m}},\beta)$, yielding an optimal price and revenue that approach $\mu$ when $m$ tends to infinity. This result also follows from Chen et al. (2022a), Giannakopoulos et al. (2023) who consider robust pricing when the mean and variance (and not the bound on support) are known.
For ease of exposition, consider $\sigma=\underaccent{\bar}{\sigma}=\bar{\sigma}$, in which case the low price is guaranteed to be optimal when $\bar{\upsilon}_{1}\geq p^{*}_{h}(\mu,\frac{\underaccent{\bar}{\sigma}}{\sqrt{m}},\beta)$, so for $m\geq m^{*}$ with $m^{*}$ the solution to
$\mu-\frac{\sigma^{2}}{m^{*}(\beta-\mu)}=\beta-\sqrt{\beta\left(\beta-\mu-\frac{\sigma^{2}}{m^{*}\mu}\right)}$ (37)
given by
$m^{*}=\left(\frac{\left(1+\sqrt{\frac{\beta+3\mu}{\beta-\mu}}\right)\beta+2\mu}{2\mu^{2}(\beta-\mu)}\right)\sigma^{2}.$ (38)
While this result is not tight, it shows that the bundle size needed to guarantee that the low price $p^{*}_{l}(\mu,\frac{\sigma}{\sqrt{m}})$ is optimal grows proportionally with the valuation variance, and that bundling in this setting with i.i.d. valuations thus leads to low pricing when the bundle is large enough.
6 Low and high pricing in service operations management
We next consider monopoly pricing for delay-prone services, a setting that is technically more involved than the classical monopoly pricing problem considered so far, but again leads to switching from low to high prices when the risk becomes large enough.
Consider a monopolistic service firm that operates as an $M/M/1$ queue. Potential customers arrive according to a Poisson process with arrival rate $\lambda$. Customers who decide to join the queue are served in order of arrival by a single server. Service times are exponentially distributed with rate $\theta$. Customers have individual service valuations that are modeled as i.i.d. random variables. In traditional models the valuation distribution is assumed to be known to the system manager, an assumption that we will relax below. Customers know their valuation but cannot observe the queue length, and decide whether to join or not only based on their assessment of the expected waiting times. Such queues are called unobservable queues and widely studied in the field of rational queueing theory; see Hassin and Haviv (2003) for an extensive overview.
For notational consistency, denote the random valuation of service by $X$, so that a customer only joins the queue when this $X$ exceeds the sum of price $p$ and holding costs $hW(\gamma)$ with $W(\gamma)$ some continuous function that increases with $\gamma$, where a standard choice would be $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$, the expected waiting time in an $M/M/1$ queue with effective arrival rate $\gamma$ and service rate $\theta$. The parameter $h$ is the cost of waiting per time unit and measures the customer’s delay sensitivity. Let $\mathbb{P}$ denote the distribution of $X$. An arbitrary customer then joins with probability $\mathbb{P}(X>p+hW(\gamma))$. (For mathematical convenience we use a strict inequality $>$ instead of $\geq$. Note that by Lemma 3.1 this will not affect the worst-case tail bound for the ambiguity sets considered in this paper.) Since customers arrive according to a Poisson process with rate $\lambda$, the effective arrival rate $\gamma=\gamma(p,\mathbb{P})$ solves
$\gamma=\lambda\mathbb{P}(X>p+hW(\gamma)).$ (39)
It is a priori not clear if (39) has a solution for given $p$ and $\mathbb{P}$; we will return to this matter later on. Note that $\lambda$ and $h$ are considered to be given constants at this point.
The above idea that individual utility leads to a global equilibrium is one of the key concepts in rational queueing theory and explained in detail in the book Hassin and Haviv (2003). See e.g. Naor (1969), Dewan and Mendelson (1990), Chen and Frank (2004), Taylor (2018), Baron et al. (2023), Chen et al. (2022b) for examples of pricing in this area, and in particular Cachon and Feldman (2011), Haviv and Randhawa (2014) consider monopoly pricing for the same model as in this section. However, all the above references assume that the valuation distribution is known, and satisfies certain regularity conditions, whereas we shall instead pursue a robust approach for when only limited information on the valuation distribution is available. Regularity conditions in Haviv and Randhawa (2014) include that the valuation distribution has a continuously differentiable density and a nondecreasing hazard rate, which warrants existence and uniqueness of the solution to (39).
A revenue-maximizing seller seeks to solve
$\sup_{p}p\cdot\gamma(p,\mathbb{P})$ (40)
and we consider the robust setting where the seller does not know the distribution $\mathbb{P}$ of $X$, but instead only knows that the valuation distribution lies in the ambiguity set $\mathcal{P}$. Taking a pessimistic approach, the seller then sets a deterministic price that maximizes the expected revenue under the worst-case scenario.
It is not guaranteed that (39) has a solution for given $p$ and $\mathbb{P}\in\mathcal{P}$. We therefore set $\gamma(p,\mathbb{P})=\infty$ if no solution to (39) exists. We assume that for every $p$, there is at least one $\mathbb{P}\in\mathcal{P}$ for which $\gamma(p,\mathbb{P})<\infty$. This is in particular true for ambiguity sets satisfying the two conditions in Theorem 43, as becomes clear from the proof. We define
$\gamma^{*}(p,\mathcal{P})=\inf_{\mathbb{P}\in\mathcal{P}}\gamma(p,\mathbb{P}),$ (41)
and then the objective of a robust seller is to solve
$\sup_{p}\,p\cdot\gamma^{*}(p,\mathcal{P}).$ (42)
Theorem 6.1.
Let $p>0$ and ambiguity set $\mathcal{P}$ be given, and assume that:
  • (i)
    $f(x)\mathrel{\mathop{:}}=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ is continuous.
  • (ii)
    $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ is attained by some distribution $\mathbb{P}_{x}$ for every $x$.
Then $\gamma^{*}=\gamma^{*}(p)$ defined in (41) is the unique solution to the equation
$\gamma^{*}=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(\gamma^{*})).$ (43)
Proof. Recall that $W$ is an increasing function and so if $x<x^{\prime}$, then also $p+hW(x)<p+hW(x^{\prime})$. This implies that $\lambda\cdot f(x)$ is non-increasing in $x$. Because $g(x)=x$ is an increasing function in $x$, it follows that the equation
$x=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ (44)
has a unique solution $x^{*}$. We will next argue that $x^{*}\geq\gamma^{*}$. Recall that $\gamma^{*}=\inf_{\mathbb{P}\in\mathbb{P}}\gamma(p,\mathbb{P})$ with $\gamma(p,\mathbb{P})$ the unique solution to the equation
$\gamma=\lambda\cdot\mathbb{P}(X>p+hW(\gamma))$ (45)
for the given distribution $\mathbb{P}$. Because of the second assumption we know that there is a distribution $\mathbb{P}_{x^{*}}$ so that
$x^{*}=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x^{*}))=\lambda\cdot\mathbb{P}_{x^{*}}(X>p+hW(x^{*})).$ (46)
This means that $\gamma^{*}\leq x^{*}<\infty$.
We will next argue that $\gamma^{*}=x^{*}$ using a proof by contradiction. Suppose that equality does not hold, i.e., $x^{*}>\gamma^{*}$. Then there exists a distribution $\mathbb{P}^{\prime}\in\mathcal{P}$ where $\gamma^{\prime}=\gamma^{\prime}(p,\mathbb{P}^{\prime})$ satisfies
$\gamma^{\prime}=\lambda\cdot\mathbb{P}^{\prime}(X>p+hW(\gamma^{\prime}))$ (47)
and $\gamma^{*}\leq\gamma^{\prime}<x^{*}$. Because the function $x\mapsto\mathbb{P}^{\prime}(X>p+hW(x))$ is non-increasing, it follows that
$\mathbb{P}^{\prime}(X>p+hW(x^{*}))\leq\mathbb{P}^{\prime}(X>p+hW(\gamma^{\prime}))=\frac{\gamma^{\prime}}{\lambda}<\frac{x^{*}}{\lambda}=\mathbb{P}_{x^{*}}(X>p+hW(x^{*})).$ (48)
By definition of $\mathbb{P}_{x^{*}}$, i.e., the second assumption of the theorem statement, this then implies that
$\mathbb{P}^{\prime}(X>p+hW(x^{*}))<\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x^{*})),$ (49)
which is a contradiction by definition of the infimum. This shows that $x^{*}=\gamma^{*}$ and completes the proof. ∎
Lemma 3.1 shows that Theorem 43 applies to all ambiguity sets considered in this paper: The function (6) is always continuous in $p$, and worst-case distributions are identified in the proof of Lemma 3.1 in Appendix A. Take as example the ambiguity set $\mathcal{P}(\mu,\beta)$ of all distributions with given mean $\mu$ and maximal valuation $\beta$. Lemma 3.1 with $\underaccent{\bar}{\sigma}=0$ and $\bar{\sigma}=\sqrt{\mu(\beta-\mu)}$ then yields
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\beta)}\mathbb{P}(X>x)=\frac{\mu-x}{\beta-x},\quad x\in(0,\mu),$ (50)
and hence, by taking $x=p+hW(\gamma^{*})$, we have that $\gamma^{*}$ solves
$\frac{\mu-p-hW(\gamma^{*})}{\beta-p-hW(\gamma^{*})}=\frac{\gamma^{*}}{\lambda}.$ (51)
For $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$ the worst-case arrival rate $\gamma^{*}$ can be obtained in closed-form. Moreover, for this case, we can also derive an explicit upper bound on $\gamma^{*}$ using the fact that $\mu-p-hW(\gamma^{*})$ needs to be nonnegative. This implies that
$\gamma^{*}\leq\gamma_{\max}\mathrel{\mathop{:}}=\theta-\frac{\theta h}{\theta\mu+h-\theta p},$ (52)
and this bound becomes exact in the large market size limit $\lambda\to\infty$. This also gives an upper bound $p\gamma_{\max}$ on the worst-case revenue, which is maximized by
$p_{\max}\mathrel{\mathop{:}}=\mu-\frac{\sqrt{h^{2}+\theta h\mu}-h}{\theta}.$ (53)
Pricing in this service system depends on the delay function $W(\cdot)$, the service rate $\theta$ and the additional parameters $h$ and $\theta$. Observe that for $h=0$ we retrieve the classical monopoly pricing problem. See Figure LABEL:plotAB for an example. Knowing the mean and maximal value thus leads to a unique optimal price that decreases with $\theta$. As in the original monopoly pricing problem with information on the mean and maximal value, there is no low-high pricing phenomenon. The result, however, is interesting in its own right, as distribution-free pricing for these service systems has not been covered in the literature.
Now assume the seller knows the mean $\mu$, bound on support $\beta$ and the standard deviation $\sigma$. Theorem 43 implies that the worst-case tail bound from Lemma 3.1 for $\underaccent{\bar}{\sigma}=\bar{\sigma}$ can be used to solve the max-min pricing (42). Figure LABEL:FigmainhQ3 shows how the arrival rate $\gamma^{*}$ depends on $\sigma$, and how this translates into a bimodal revenue, as in the original monopoly pricing problem. Again, there is a switch from low to high pricing as the variance grows beyond a certain threshold.
Figures LABEL:figmain1hQ and LABEL:figmain1hoptQ show that increasing $h$ or decreasing $\beta$ also forces switching from low to high pricing. Penalizing waiting more heavily indeed makes low pricing with high demand less profitable. And why is it that restricting valuations leads the seller to post high prices and vice versa? Allowing larger valuations increases the set of valuation distributions, providing more opportunity to concentrate mass on low values while meeting the mean constraint. This leaves the seller no other option than selecting a low price. Similarly, preventing the adversary choosing the worst-case distribution from assigning high valuations forces the seller to post a high price. These are universal insights that hold beyond the examples of monopoly pricing and delay-prone services considered in this paper.
7 Conclusions and outlook
Transitioning from low to high pricing in response to increased market volatility has not been previously documented in the existing literature on robust pricing. When next to the mean, the seller is only aware of the standard deviation, this leads to relatively lower prices that decrease with risk. Similarly, relying exclusively on valuation bounds results in risk-insensitive pricing. The distinctive switch from low to high pricing occurs only when the seller possesses, in addition to the mean and valuation bounds, sufficiently precise information about the standard deviation. This specific combination of factors contributing to the phenomenon might explain why it has been overseen so far. Another reason might be the dominance of low pricing strategies in prior research, coupled with a focus on the intuition that prices should decrease with the level of risk or volatility.
As illustrated in the example of Beta-distributed valuations discussed in the introduction, instances of large or extreme volatility translate into worst-case scenarios, akin to all-or-nothing scenarios. When these scenarios become sufficiently likely, the seller shifts strategies from low pricing and aggressive market penetration to high prices tailored exclusively for those users with the highest valuations. This strategic shift emphasizes the imperative to tailor pricing strategies to accommodate different levels of market uncertainty, particularly in instances of substantial market volatility. We anticipate that the transition to higher pricing due to increased risk, as elucidated in this study via a basic yet parsimonious pricing model, holds broader applicability in more generalized or detailed models and real-world settings. We have demonstrated this already for the setting of delay-prone services.
Exploring the occurrence of low-high pricing in robust settings with partial demand information captured in terms of ambiguity sets is worthwhile to investigate in more detail. Ambiguity sets not explored in this paper include those conditioned on the first three or more moments, or distance-based ambiguity sets centered around a reference distribution, with distance measured in terms of metrics like $\phi$-divergence, Kullback-Leibler distance and Wasserstein distance. While necessitating a distinct mathematical approach, we anticipate that these alternative ambiguity sets may also yield instances of low-high pricing, particularly for sufficiently high levels of ambiguity.
Acknowledgements. This research was supported by NWO Vici grant 202.068. We also thank Alex Suzdaltsev for exchanging ideas about some unpublished results in Suzdaltsev (2018).
Appendix A Proof of Lemma 3.1
Let $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$ with $\mathcal{B}$ the Borel sigma-algebra on $\mathbb{R_{+}}$. We will solve $\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p)$, which can be rewritten as
$\!\inf_{\mathbb{P}\in\mathcal{M}}$ $\int_{x}\mathds{1}{\{x>p\}}{\rm d}\mathbb{P}(x),$ (54)
s.t. $\int_{x}{\rm d}\mathbb{P}(x)=1,\ \int_{x}x{\rm d}\mathbb{P}(x)=\mu,\ \int_{x}x^{2}{\rm d}\mathbb{P}(x)\geq\underaccent{\bar}{\sigma}^{2}+\mu^{2},\ \int_{x}x^{2}{\rm d}\mathbb{P}(x)\leq\bar{\sigma}^{2}+\mu^{2}.$
Consider the corresponding dual,
$\sup_{\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}}$ $\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2}),$ (55)
s.t. $\mathds{1}{\{x>p\}}\geq\lambda_{0}+\lambda_{1}x+(\lambda_{2}+\lambda_{3})x^{2}=:F(x),\ \forall x\in[0,\beta],\lambda_{2}\geq 0,\lambda_{3}\leq 0.$
The proof is structured in the following manner: Initially, four scenarios are considered, each corresponding to a feasible solution for the dual problem. Subsequently, we construct four primal solutions corresponding to these scenarios. The final step is the demonstration that strong duality is maintained across all possible values of $p$, with each scenario uniquely aligned with a specific interval of $p$. We now consider these scenarios.
In the first scenario, $F(x)$ intersects $\mathds{1}\{x>p\}$ in $\{p,\alpha\}$ with $\alpha\in[p,\beta]$. In the second scenario, the intersection occurs in the points $\{p,\beta\}$. For the third scenario, the points of intersection are identified at $\{0,p,\beta\}$. Finally, in the fourth scenario, the intersection occurs at the interval $[0,p]$. The four scenarios are depicted in Figure 7.
[…]
Figure 7: All four scenarios with $\mu=1$, $\underset{\bar{}}{\sigma}=0.1$, $\bar{\sigma}=1$, and $\beta=4$.
We now present four viable dual solutions that correspond to these scenarios. For scenario 1, the conditions $F(p)=0$, $F(\alpha)=1$ and $F^{\prime}(\alpha)=0$ are observed. Moreover, this scenario has $F$ as a downward-sloping parabola. To align with these conditions, we set $\lambda_{2}=0$, as $\lambda_{3}\leq 0$. This gives
$\lambda_{0}=-\frac{2\alpha p-p^{2}}{(\alpha-p)^{2}},\ \lambda_{1}=\frac{2\alpha}{(\alpha-p)^{2}},\ \ \lambda_{2}=0,\ \lambda_{3}=-\frac{1}{(\alpha-p)^{2}}.$ (56)
The point $\alpha$ can be determined by considering a two-point distribution with support $\{p,\alpha\}$, mean $\mu$, and variance $\bar{\sigma}^{2}$, yielding
$\alpha=\mu+\frac{\bar{\sigma}^{2}}{\mu-p}.$
Consequently,
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{(\mu-p)^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}}.$ (57)
In scenario 2, we observe that $F(p)=0$ and $F(\beta)=1$. In addition, $F$ is characterized as a line, leading $\lambda_{2}=\lambda_{3}=0$. This gives dual solution
$\lambda_{0}=-\frac{p}{\beta-p},\ \lambda_{1}=\frac{1}{(\beta-p)},\ \lambda_{2}=0,\ \lambda_{3}=0,$ (58)
and
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{\mu-p}{\beta-p}.$ (59)
In scenario 3 we observe that $F(0)=0$, $F(p)=0$, and $F(\beta)=1$. Moreover, $F$ is an upward-sloping parabola. Hence, in line with the functional form, we set $\lambda_{3}=0$, as $\lambda_{2}\geq 0$. This results in
$\lambda_{0}=0,\ \lambda_{1}=-\frac{p}{\beta(\beta-p)},\ \lambda_{2}=\frac{1}{\beta(\beta-p)},\ \lambda_{3}=0,$ (60)
and
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (61)
Finally, in scenario 4, we note that $F(x)=0$, for all $x$ in the interval $[0,\beta]$. This gives $\lambda_{0}=0$, $\lambda_{1}=0$, $\lambda_{2}=0$, and $\lambda_{3}=0$. Hence,
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=0.$ (62)
Subsequently, we formulate four primal solutions, each with the same objective value as those obtained from the dual problem solutions. This equality in objective values asserts the claim of strong duality. A constructive method for deriving these primal solutions is now presented. Assuming strong duality holds, we have
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}^{*}(x)=\int_{x}(\lambda^{*}_{0}+\lambda^{*}_{1}\mu+\lambda^{*}_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda^{*}_{3}(\bar{\sigma}^{2}+\mu^{2})){\rm d}\mathbb{P}^{*}(x).$ (63)
Furthermore, due to dual feasibility, we obtain
$\lambda^{*}_{0}+\lambda^{*}_{1}\mu+\lambda^{*}_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda^{*}_{3}(\bar{\sigma}^{2}+\mu^{2})-\mathds{1}\{x>p\}\geq 0,$ (64)
holding pointwise for each $x\in[0,\beta]$. Equation (63) in conjunction with equation (64) is known as the complementary slackness relationship in semi-infinite linear programming. A direct implication of this principle is that the atoms of the optimal primal solution coincide with the intersection of $\lambda^{*}_{0}+\lambda^{*}_{1}x+(\lambda^{*}_{2}+\lambda^{*}_{3})x^{2}$ and the indicator function $\mathds{1}\{x>p\}$. Following this, we will present four viable dual solutions corresponding to these scenarios.
In scenario 1, we construct a two-point distribution with atoms $\{p,\alpha\}$. The corresponding probabilities are
$p_{p}=\frac{\bar{\sigma}^{2}}{\mu^{2}+\bar{\sigma}^{2}+p^{2}-2\mu p},\quad p_{\alpha}=1-p_{p},$ (65)
leading to
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{\alpha}=\frac{(\mu-p)^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}}.$ (66)
However, this distribution’s feasibility requires $\alpha\leq\beta$, necessitating $p\in[0,\bar{\upsilon}_{1}]$.
For scenario 2, the construction involves a two-point distribution with atoms $\{p,\beta\}$. The probabilities are
$p_{p}=\frac{\beta-\mu}{\beta-p},\quad p_{L}=\frac{\mu-p}{\beta-p},$ (67)
resulting in
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{L}=\frac{\mu-p}{\beta-p}.$ (68)
However, the feasibility criterion necessitates
$\underaccent{\bar}{\sigma}^{2}\leq(p-\mu)^{2}p_{p}+(\beta-\mu)^{2}p_{L}\leq\bar{\sigma}^{2},$ (69)
requiring $p\in[\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1}]$.
For scenario 3, we consider a three-point distribution with probability masses on the points $\{0,p,\beta\}$ and corresponding probabilities
$p_{0}=\frac{(\beta-\mu)(p-\mu)+\underaccent{\bar}{\sigma}^{2}}{Lp},\quad p_{p}=1-p_{0}-p_{L},\quad p_{L}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (70)
This yields
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{L}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (71)
However, this distribution is only feasible when $p_{0}\geq 0$ and $p_{L}\geq 0$, which requires $p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$.
Scenario 4 sees a two-point distribution with variance $\underaccent{\bar}{\sigma}^{2}$ and atoms $\{\bar{\alpha},p\}$ with $\bar{\alpha}=\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{p-\mu}$. The corresponding probabilities are
$p_{\bar{\alpha}}=1-p_{p},\quad p_{p}=\frac{(p-\mu)^{2}}{(p-\mu)^{2}+\underaccent{\bar}{\sigma}^{2}},$ (72)
and hence
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=0.$ (73)
However, this distribution is only feasible when $\beta\geq 0$, requiring $p\in[\underaccent{\bar}{\upsilon}_{2},\beta]$. Since the primal and dual objective values are the same, we know that strong duality holds. Therefore, these objective values are optimal. In addition, the objective value is left-continuous in $p$. Therefore, we can apply Lemma A.1, stated below, and conclude that
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p).$ (74)
This completes the proof.
Lemma A.1.
Let $p\in(0,\beta]$, $\mathcal{P}$ any ambiguity set, and assume that
$\lim_{\epsilon\xrightarrow{}p^{-}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>\epsilon)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (75)
Then
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (76)
Proof. Let $p\in(0,\beta]$ and $\epsilon>0$. The following identity trivially holds:
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon).$ (77)
However, since $\epsilon$ is chosen arbitrarily, we also have that
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\lim_{\epsilon\xrightarrow{}0^{+}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon)=\lim_{\epsilon\xrightarrow{}p^{-}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>\epsilon)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (78)
Hence, we obtain
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p),$ (79)
which proves the claim. ∎
Appendix B Remaining proofs of pricing results
B.1 Proof of Lemma 4.2
We first set $\mu=1$ without loss of generality. Recall $\hat{R}(p)$ from (18) and define $\hat{R}(p;\mu,\sigma,\beta)\mathrel{\mathop{:}}=\hat{R}(p)$. In addition, consider the scaled parameters $\nu=\frac{\sigma}{\mu}$, $\tau=\frac{\beta}{\mu}$, and $\psi=\frac{p}{\mu}$. From (18) it follows that
$\hat{R}(p;\mu,\sigma,\beta)=\hat{R}(\psi\mu;\mu,\nu\mu,\tau\mu)=\mu\cdot\hat{R}(\psi;1,\nu,\tau).$ (80)
Hence, using the substitution $p=\psi\mu$, we have
$\arg\sup_{p}\hat{R}(p;\mu,\sigma,\beta)=\mu\arg\sup_{\psi}\hat{R}(\psi;1,\nu,\tau).$ (81)
We will now proceed with $\mu=1$. Consider that for $\underaccent{\bar}{\upsilon}_{1}$ to be the unique global optimum, the following ordering needs to hold:
$p^{*}_{h}(\sigma,\beta)<\underaccent{\bar}{\upsilon}_{1}<p^{*}_{l}(\sigma).$ (82)
We will show that this ordering never occurs by considering two cases that exhaust all possibilities. The first case considers $\sigma\geq 1$. We claim the following:
$p^{*}_{h}(\sigma,\beta)\geq 1\geq\underaccent{\bar}{\upsilon}_{1}.$ (83)
The second inequality is trivial. To see why the first inequality holds true, note that
$\beta-\sqrt{\beta(\beta-(1+\sigma^{2}))}\geq 1$ $\Leftrightarrow\beta-1\geq\sqrt{\beta^{2}-\beta-\beta\sigma^{2}}$
$\Leftrightarrow\beta^{2}-2\beta+1\geq\beta^{2}-\beta-\beta\sigma^{2}$
$\Leftrightarrow 1+\beta\sigma^{2}\geq\beta.$ (84)
This is true since both $\sigma^{2}\geq 1$ and $\beta\geq 1$.
The second case considers $\sigma\in(0,1)$. Denote
$\beta_{l}\mathrel{\mathop{:}}=\frac{1}{2}\left(2+\sigma^{2}+A^{\frac{1}{3}}+\frac{\sigma^{4}}{A^{\frac{1}{3}}}\right)\text{with }A\mathrel{\mathop{:}}=\sigma^{4}(\sqrt{1+\sigma^{2}}+1)^{2},$ (85)
as the unique value of $\beta$ for which $p^{*}_{l}(\sigma)=\underaccent{\bar}{\upsilon}_{1}$, and
$\beta_{h}\mathrel{\mathop{:}}=\frac{2-2^{\frac{4}{3}}\sigma^{4}B^{-\frac{1}{3}}-2^{\frac{2}{3}}B^{\frac{1}{3}}}{2(1-\sigma^{2})}\text{with }B\mathrel{\mathop{:}}=-2\sigma^{8},$ (86)
as the unique value for which $p^{*}_{h}(\sigma,\beta)=\underaccent{\bar}{\upsilon}_{1}$. For (82) to hold, it has to be the case that $\beta_{h}<\beta_{l}$. Hence, we will show that the opposite, namely $\beta_{h}\geq\beta_{l}$ is always true. It is sufficient to demonstrate that $\beta_{h}\geq(1-\sigma^{2})\beta_{h}\geq\beta_{l}$. This results in the following inequality:
$2\sigma^{\frac{4}{3}}+2\sigma^{\frac{8}{3}}\geq\sigma^{2}+\sigma^{\frac{4}{3}}(\sqrt{1+\sigma^{2}}+1)^{\frac{2}{3}}+\frac{\sigma^{\frac{8}{3}}}{(\sqrt{1+\sigma^{2}}+1)^{\frac{2}{3}}}.$ (87)
Dividing by $\sigma^{\frac{4}{3}}$ and substituting $c$ for $\sigma^{\frac{2}{3}}$ results in
$2+2c^{2}\geq c+(\sqrt{1+c^{3}}+1)^{\frac{2}{3}}+\frac{c^{2}}{(\sqrt{1+c^{3}}+1)^{\frac{2}{3}}},$ (88)
which is the same as
$2+2c^{2}-c\geq(\sqrt{1+c^{3}}+1)^{-\frac{2}{3}}\left[(\sqrt{1+c^{3}}+1)^{\frac{4}{3}}+c^{2}\right].$ (89)
Since $\sqrt{1+c^{3}}+1\geq 2$ for every $c\in(0,1)$, we have
$\frac{2}{3}\geq 2^{-\frac{2}{3}}\geq(\sqrt{1+c^{3}}+1)^{-\frac{2}{3}}.$ (90)
Hence, combining (89) with (90) results in
$2+2c^{2}-c\geq\frac{2}{3}\left[(\sqrt{1+c^{3}}+1)^{\frac{4}{3}}+c^{2}\right],$ (91)
which is equivalent to
$3+2c^{2}-\frac{3}{2}c\geq(\sqrt{1+c^{3}}+1)^{\frac{4}{3}}.$ (92)
Denote the left-hand side as $r(c)\mathrel{\mathop{:}}=3+2c^{2}-\frac{3}{2}c$ and the right-hand side as $t(c)\mathrel{\mathop{:}}=(\sqrt{1+c^{3}}+1)^{\frac{4}{3}}$. Note that the quadratic function $r(c)$ has one local minimum at $c^{*}=\frac{3}{8}$ and that $t(c)$ is strictly increasing in $c$. By inspection, one can verify that (92) is true for both $c=0$ and $c=c^{*}$. Since $r(c)$ is decreasing on the interval $(0,c^{*}]$, and $t(c)$ increasing, it follows that (92) is true for all $c\in(0,c^{*}]$. Both functions are increasing on the interval $(c^{*},1]$. We will argue that it suffices to check (92) for a finite number of points to conclude that it holds for all $c\in(c^{*},1)$.
Assume that for a given $\epsilon>0$ and $d,d+\epsilon\in(c^{*},1]$, we have $r(d)>t(d+\epsilon)$. Then (92) holds for any $c\in[d,d+\epsilon]$, as
$r(c)\geq r(d)\geq t(d+\epsilon)\geq t(c),$ (93)
since $r$ and $t$ are both continuous and increasing. Therefore, we can make a discretization of the interval $(c^{*},1]$ formed by $K$ points $c^{*}=c_{0}<c_{1}<...<c_{K}=1$ with $c_{k}=c_{0}+\epsilon*k$, for all $k\in\{1,...,K\}$. As it turns out, $K=6$ is the smallest number for which this method works and has corresponding $\epsilon=\frac{1-c^{*}}{6}=\frac{5}{48}$. Table 2 shows the function evaluations that together prove (92).
$c_{k}$ $r(c_{k})\approx$ $t(c_{k}+\epsilon)\approx$
$c_{0}=\frac{3}{8}$ 2.72 2.61
$c_{1}=\frac{23}{48}$ 2.74 2.68
$c_{2}=\frac{7}{12}$ 2.81 2.78
$c_{3}=\frac{11}{16}$ 2.91 2.90
$c_{4}=\frac{19}{24}$ 3.07 3.06
$c_{5}=\frac{43}{48}$ 3.26 3.24
Table 2: Numerical calculations required in the proof of (92).
This proves that (82) never occurs when $\sigma\in(0,1)$. Hence, $\underaccent{\bar}{\upsilon}_{1}$ is never the unique global optimum.
B.2 Proof of Corollary 5.2
With shorthand notation $r\mathrel{\mathop{:}}=\frac{\mu}{\sigma}$ and $r_{+}\mathrel{\mathop{:}}=\sqrt{1+(\frac{\mu}{\sigma})^{2}}>r$, observe that
$\frac{\partial\rho_{l}(\mu,\bar{\sigma})}{\partial r}=\dfrac{\left(\sqrt[3]{r+r_{+}}+\sqrt[3]{r-r_{+}}\right)^{2}\left(\left(r+r_{+}\right)^{\frac{2}{3}}-\left(r-r_{+}\right)^{\frac{2}{3}}\right)}{2r^{2}r_{+}\left(r-r_{+}\right)^{\frac{2}{3}}\left(r+r_{+}\right)^{\frac{2}{3}}}\geq 0,$ (94)
as $r>0$ and
$\left(r+r_{+}\right)^{\frac{2}{3}}-\left(r-r_{+}\right)^{\frac{2}{3}}\geq 0.$ (95)
Hence, $\rho_{l}(\mu,\bar{\sigma})$ increases in $\mu$ and decreases in $\bar{\sigma}$. Next, denote $s\mathrel{\mathop{:}}=\frac{\beta}{\mu}$ and observe that
$\frac{\partial\rho_{m}(\mu,\beta)}{\partial s}=2\left(1-\dfrac{2s-1}{2\sqrt{\left(s-1\right)s}}\right)\leq 0.$ (96)
Therefore, $\rho_{m}(\mu,\beta)$ increases in $\mu$ and decreases in $\beta$. With $q\mathrel{\mathop{:}}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}$ observe that
$\frac{\partial\rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)}{\partial q}=\dfrac{1}{\sqrt{1-q}}-1\geq 0,$ (97)
since $q=\frac{\underaccent{\bar}{\upsilon}_{2}}{\beta}<1$. This shows that $\rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)$ increases in $\underaccent{\bar}{\sigma}$ and decreases in $\beta$. Next, recall that $p^{*}_{l}(\bar{\sigma})$ is the unique real interior solution of $R_{1}(p)$, so solving
$(p^{*}_{l}(\bar{\sigma}))^{3}-3\mu(p^{*}_{l}(\bar{\sigma}))^{2}+3(\mu^{2}+\bar{\sigma}^{2})p^{*}_{l}(\bar{\sigma})=\mu(\mu^{2}+\bar{\sigma}^{2}).$ (98)
To simplify the analysis, we change variables by substituting $\mu_{2}\mathrel{\mathop{:}}=\mu^{2}+\bar{\sigma}^{2}$ (the maximal second moment), and write (98) as
$(p^{*}_{l}(\mu_{2}-\mu))^{3}-3\mu(p^{*}_{l}(\mu_{2}-\mu))^{2}+3\mu_{2}p^{*}_{l}(\mu_{2}-\mu)=\mu\mu_{2}.$ (99)
By keeping $p^{*}_{l}(\mu_{2}-\mu)$ implicit while isolating $\mu$ and $\mu_{2}$ in (99), we can write (98) as
$L_{1}(p^{*}_{l}(\mu_{2}-\mu),\mu_{2})\mathrel{\mathop{:}}=\frac{(p^{*}_{l}(\mu_{2}-\mu))^{3}+3p^{*}_{l}(\mu_{2}-\mu)\mu_{2}}{3(p^{*}_{l}(\mu_{2}-\mu))^{2}+\mu_{2}}=\mu,$ (100)
or
$L_{2}(p^{*}_{l}(\mu_{2}-\mu),\mu)\mathrel{\mathop{:}}=\frac{-(p^{*}_{l}(\mu_{2}-\mu))^{2}(p^{*}_{l}(\mu_{2}-\mu)-3\mu)}{3p^{*}_{l}(\mu_{2}-\mu)-\mu}=\mu_{2}.$ (101)
Differentiating both sides of (100) with respect to $\mu$ yields
$\frac{\partial L_{1}(p^{*}_{l}(\mu_{2}-\mu),\mu_{2})}{\partial p^{*}_{l}(\mu_{2}-\mu)}\cdot\frac{\partial p^{*}_{l}(\mu_{2}-\mu)}{\partial\mu}=\frac{3((p^{*}_{l}(\mu_{2}-\mu)))^{2}-\mu_{2})^{2}}{(3p^{*}_{l}(\mu_{2}-\mu)^{2}+\mu_{2})^{2}}\cdot\frac{\partial p^{*}_{l}(\mu_{2}-\mu)}{\partial\mu}=1,$ (102)
showing that $p^{*}_{l}(\bar{\sigma})$ is (strictly) increasing in $\mu$. Differentiating both sides of (101) with respect to $\mu_{2}$ gives
$\frac{\partial L_{2}(p^{*}_{l}(\mu_{2}-\mu),\mu)}{\partial p^{*}_{l}(\mu_{2}-\mu)}\cdot\frac{\partial p^{*}_{l}(\mu_{2}-\mu)}{\partial\mu_{2}}=\frac{-6p^{*}_{l}(\mu_{2}-\mu)(p^{*}_{l}(\mu_{2}-\mu)-\mu)^{2}}{(3p^{*}_{l}(\mu_{2}-\mu)-\mu)^{2}}\cdot\frac{\partial p^{*}_{l}(\mu_{2}-\mu)}{\partial\mu_{2}}=1,$ (103)
showing that $p^{*}_{l}(\bar{\sigma})$ is (strictly) decreasing in $\bar{\sigma}$. To see why $p^{*}_{m}(\beta)$ is increasing in $\mu$ and decreasing in $\beta$, we examine $\rho_{m}(\mu,\beta)={2p^{*}_{m}(\beta)}/{\mu}-1$, which is clearly increasing in $p^{*}_{m}(\beta)$. Hence, the same monotonicity properties directly apply to $p^{*}_{m}(\beta)$. Finally, $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)$ increases in $\underaccent{\bar}{\sigma}$ and decreases in $\beta$, due to $p^{*}_{h}(\underaccent{\bar}{\sigma},\beta)=p^{*}_{m}(\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu},\beta)$ and monotonicity of both $\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu}$ and $p^{*}_{m}(\beta)$.